OpenStudy (anonymous):

which of the following equations would graph a line parallel to 3y=2x-x+5

OpenStudy (anonymous):

a. 3y=x+1 b.y=x+1 c.y=2/3x+1 d.3/2y=x+1

OpenStudy (anonymous):

try to set it into a slope intercept form y= mx+b.. m being the slope, and which ever has the same slope to the equation is the parallel to the original equation

OpenStudy (anonymous):

or you can just combine the like terms from 3y=2x-x+5 then, you can solve it like that for this case

OpenStudy (anonymous):

I'm lost

OpenStudy (anonymous):

like term for 3y=2x-x+5 is 3y=x+5, from the giving choices, which one has 3y = x?

OpenStudy (anonymous):

a

OpenStudy (anonymous):

correct, and that is because 3y=2x-x+5 and 3y=x+1 has the same slope when you put it in slope intercept form

OpenStudy (anonymous):